OpenStudy (anonymous):

I have two functions I cannot figure out. 1) f(-1)=2*-1^2 + -1 2)f(-1)=|-1+3| help!!

OpenStudy (anonymous):

pemdas parenthesis exponent multiplication division addition subtraction

OpenStudy (anonymous):

1) f(-1) = 2* (1) + (-1) = 2 - 1 = 1 2) f(-1) = I2I = 2

OpenStudy (anonymous):

absolute value signs act like parenthesis

OpenStudy (anonymous):

Im not sure I understand? but ty very much

OpenStudy (anonymous):

order of operations you do everything in the parenthesis first then exponents then multiplication etc etc etc

OpenStudy (anonymous):

start by looking at your first problem 2*-1^2 + -1 no parenthesis, so we go to the exponent -1^2= 1 2*1 + -1 now we go to the multiplication 2*1=2 2+ -1 then we add whats left 2+-1 =1

OpenStudy (anonymous):

I understand the PEMDAS concept...so for f(3) =2 (3)^2+3 would would you get, I got 21

OpenStudy (anonymous):

what would you get is what I meant to say. :)

OpenStudy (anonymous):

i got 21 2*9+3 18+3 21